A survey was taken to determine the proportion of recent BCIT Marketing graduates who achieved a starting salary of at least $50,000 per year.

A random sample of 100 graduates resulted in a confidence interval of 12% ≤Π≤ 36%.

What is the confidence level used for this estimate of the population proportion? Use z, not t

Give the confidence level as a percent (without the % symbol), rounded to the nearest whole percent:
Answer #1

Solution:

n = 100

Given a confidence interval for p is

12% ≤ Π ≤ 36%

0.12  ≤ Π ≤ 0.36

Upper limit = 0.36

Lower limit = 0.12

\hat p = (Upper + Lower)/2 = (0.36+0.12)/2 = 0.24

Margin of error = E = (Upper - Lower)/2 = (0.36-0.12)/2 = 0.12

But ,

E = 2012 *  Vp(1-P)/n

\therefore 0.12 =  2012 *   \sqrt{}[0.24 *(1 - 0.24)/100]

\therefore 0.12 =   2012 * 0.042708313

\therefore2012 =  2.81

Use z table .Go to column of z . See 2.8 in this column and then see corresponding probability in the row at .01

It is 0.9975

So ,  1- \alpha /2 = 0.9975

\alpha/2 = 0.0025

\alpha = 0.005

Now ,

confidence level = 1 - \alpha = 1 - 0.005 = 0.995 =99.5% = 99%

Answer : 99